# APCodec: A Neural Audio Codec with Parallel Amplitude and Phase Spectrum Encoding and Decoding

### v1

@@ Abstract
This paper introduces a novel neural audio codec targeting high waveform sampling rates and low bitrates named APCodec, which seamlessly integrates the strengths of parametric codecs and waveform codecs. The APCodec revolutionizes the process of audio encoding and decoding by concurrently handling the amplitude and phase spectra as audio parametric characteristics like parametric codecs. It is composed of an encoder and a decoder with the modified ConvNeXt v2 network as the backbone, connected by a quantizer based on the residual vector quantization (RVQ) mechanism. The encoder compresses the audio amplitude and phase spectra in parallel, amalgamating them into a continuous latent code at a reduced temporal resolution. This code is subsequently quantized by the quantizer. Ultimately, the decoder reconstructs the audio amplitude and phase spectra in parallel, and the decoded waveform is obtained by inverse short-time Fourier transform. To ensure the fidelity of decoded audio like waveform codecs, spectral-level loss, quantization loss, and generative adversarial network (GAN) based loss are collectively employed for training the APCodec. To support low-latency streamable inference, we employ feed-forward layers and causal convolutional layers in APCodec, incorporating a knowledge distillation training strategy to enhance the quality of decoded audio. Experimental results confirm that our proposed APCodec can encode 48 kHz audio at bitrate of just 6 kbps, with no significant degradation in the quality of the decoded audio. At the same bitrate, our proposed APCodec also demonstrates superior decoded audio quality and faster generation speed compared to well-known codecs, such as SoundStream, Encodec, HiFi-Codec and AudioDec.
@@ I Introduction
The waveform codecs aim at encoding the input audio waveform directly and reproducing a faithful reconstruction of the input audio waveform, such as pulse code modulation (PCM) [20]. Although waveform codecs can decode high-quality audio, they also require a higher bitrate, which increases storage and transmission costs. In recent times, end-to-end neural waveform codecs with raw waveform I/O have surfaced, offering a partial equilibrium between decoded audio quality and bitrate [21, 22, 23, 24, 25, 26]. For example, SoundStream [25] and Encodec [26] employed the residual vector quantization (RVQ) mechanism [27] to reduce the bitrate, while utilizing the losses from the HiFi-GAN vocoder [28] to ensure the fidelity of the decoded audio. Recently, researchers have made improvements addressing the issues present in current end-to-end audio codecs, primarily focusing on quantization strategies. On the one hand, in applications such as audio communication, audio codecs have incorporated variations of RVQ to decrease bitrates and improve communication efficiency [29, 30, 31]. On the other hand, in downstream tasks such as TTS, efforts have been undertaken to introduce or disentangle semantic information during the quantization stage, aiming to tailor the approach to the specific tasks [7, 8, 9]. Moreover, there have been endeavors to improve the model structure [30] or incorporate additional signal processing techniques in codecs [32]. Although these codecs have indeed enhanced the decoded audio quality and decreased bitrates, they still required more than a hundred times the downsampling and upsampling operations due to the direct waveform encoding and decoding, leading to high model complexity. Besides, direct waveform encoding and decoding could also potentially result in low generation efficiency. Some recent works have also overlooked considerations for low latency, making it challenging to achieve streamable inference [29, 30].
@@
Beyond the aforementioned challenges, encompassing bitrate, generation speed, and latency in existing audio codecs, there is presently scant research devoted to audio codecs tailored for higher waveform sampling rates (e.g., 48 kHz). Currently, neural audio codecs (e.g., SoundStream and HiFi-Codec) are mostly designed for processing audio at sampling rates of 16 kHz or 24 kHz. This limitation hinders the utilization of audio codecs for compressing high-sampling-rate audio data and poses challenges for downstream tasks like TTS, which aim to meet the demand for higher quality speech generation. The aforementioned MDCT-based parametric codecs [18, 19], while targeted at 48 kHz audio, demands an excessively high bitrate. Although AudioDec [33] can achieve 48 kHz audio codec at bitrate of 12.8 kbps, it still requires the integration of a neural vocoder and the adoption of a multi-stage training strategy, as reported in [33].
@@
To address the aforementioned challenges, this paper proposes a novel neural audio codec named APCodec. It endeavors to provide high-quality decoded audio, while maintaining a low bitrate, fast generation speed, and low latency, specifically tailored for 48 kHz audio. Like parametric codecs, the proposed APCodec regards amplitude and phase spectra as audio parameter characteristics during the encoding and decoding processes, rather than directly processing the raw waveform. A notable advantage of this approach lies in its simplicity, as it only requires uncomplicated downsampling to obtain latent codes at an appropriately low sampling rate, thereby effectively reducing the bitrate. RVQ [27] is also utilized for code quantization to further reduce the bitrate. With the objective of achieving faithful waveform reconstruction akin to waveform codecs, a comprehensive combination of spectral-level loss, quantization loss and generative adversarial network (GAN) based loss are employed to train the APCodec. To attain streamable inference, a low-latency implementation is achieved by integrating feed-forward layers and causal convolutional layers, complemented by the application of a knowledge distillation training strategy. The resulting fixed latency is only 6.67 ms for the 48 kHz audio codec. Experimental results have confirmed that the proposed APCodec can achieve high-quality 48 kHz audio codec at a bitrate of only 6 kbps with only 8$\times$ downsampling/upsampling. At the same bitrate, our proposed APCodec significantly outperforms several well-known neural codecs, e.g., SoundStream [25], Encodec [26], HiFi-Codec [29] and AudioDec [33] in terms of decoded audio quality. The APCodec also demonstrates the fastest generation speed, attaining an impressive 89$\times$ real-time performance on GPU and 5.8$\times$ real-time performance on CPU. This remarkable acceleration is attributed to its comprehensive all-frame-level processing.
@@
This paper is organized as follows: In Section II, we briefly review the advanced neural audio codecs. In Section III, we provide details on our proposed APCodec. In Section IV, we present our experimental results. Finally, we give conclusions in Section V.
II Related Works
The focus of this paper is the neural audio codec. SoundStream [25] and Encodec [26] are earlier well-known neural audio codec baseline models that directly encode and decode audio waveforms. They consist of an encoder, a quantizer, and a decoder. The encoder and decoder are designed based on causal convolutional layers for low-latency streamable inference, while the quantizer employs RVQ to reduce the bitrate. GAN-based loss, inspired by HiFi-GAN [28], along with reconstruction loss, quantization loss, and other components, is utilized to train the encoder, quantizer, and decoder jointly.
Recently, researchers have proposed several improvement methods around the framework of SoundStream and Encodec. Regarding quantization strategy, HiFi-Codec [29] has introduced group RVQ (GRVQ) to reduce information redundancy in RVQ. This allows for high-quality audio codec with less codebooks, resulting in a reduced bitrate. CBRC [30] proposes group-wise and beam-search RVQ (GB-RVQ) to mitigate quantization errors, thereby enhancing the decoded audio quality. Several latest speech codecs such as SpeechTokenizer [7], RepCodec [8] and TiCodec [9], also attempt to introduce or disentangle semantic information within the quantized discrete codes, making them better suited for downstream tasks such as TTS. In terms of model architecture, CBRC [30] introduces convolutional and bidirectional recurrent neural network (CBRN) blocks to replace conventional convolutional layers, aiming to enhance the quality of decoded audio. Additionally, some efforts leverage other tools to achieve improvements in neural audio codecs [17, 32]. For example, CQNV [17] initially encodes and quantizes 8 kHz speech, and finally restores 16 kHz speech directly through a HiFi-GAN vocoder, implying a process of bandwidth extension (BWE). This method of encoding low-sampling-rate speech contributes to reducing the bitrate.
The aforementioned neural audio codecs (e.g., SoundStream and HiFi-Codec) are designed to handle 16 kHz or 24 kHz audio. For high-sampling-rate audio codecs, there have been efforts to perform encoding, quantization, and decoding in the MDCT domain [18, 19]. However, the required bitrate is very high, thus limiting the compression efficiency. At present, there is still a noticeable deficiency in the field of audio codec within other spectral domains (e.g., amplitude and phase spectra). The recently proposed AudioDec [33] has been reported to achieve a 48 kHz audio codec at 12.8 kbps. However, AudioDec employs a sophisticated two-stage training mode, and experimental findings reveal that achieving high-quality audio decoding still necessitates the support of the HiFi-GAN vocoder.
Figure 1: Details of the model structure of the proposed APCodec. Here, Conv1D, DeConv1D, Concat, $\Phi$, STFT and ISTFT represent the 1D convolutional layer, 1D deconvolutional layer, concatenation, phase calculation formula, short-time Fourier transform and inverse short-time Fourier transform, respectively.
III Proposed Methods
Unlike some well-known neural waveform codecs (e.g., SoundStream [25] and Encodec [26]), our proposed APCodec encodes and quantizes amplitude and phase spectra extracted from the audio waveform through short-time Fourier transform (STFT). Finally, it decodes the quantized codes into amplitude and phase spectra and restores the audio waveform through inverse STFT (ISTFT). Subsequently, we will present a detailed overview of the model structure and training criteria of the proposed APCodec. Additionally, we will discuss the low-latency implementation for APCodec.
III-A Model Structure
@@ III-A1 Encoder
As illustrated in Fig. 1, the encoder takes the log amplitude spectrum $\bm{A}\in\mathbb{R}^{F\times N}$ and phase spectrum $\bm{P}\in\mathbb{R}^{F\times N}$ extracted from the audio waveform $\bm{x}\in\mathbb{R}^{T}$ using STFT as inputs and encodes them into a continuous latent code $\bm{C}\in\mathbb{R}^{F_{c}\times N_{c}}$ that contains fused amplitude and phase information in parallel. Here, $T$ represents the number of time-domain waveform samples, and $F$ and $N$ respectively represent the number of spectral frames and frequency bins. Assuming the sampling rate of $\bm{x}$ is $f_{s}$ and the frame shift of the STFT is $w_{s}$, the resulting sampling rate for the extracted amplitude and phase spectra is $f_{s}/w_{s}$, and it holds that $T=F\cdot w_{s}$. $F_{c}$ and $N_{c}$ denote the number of frames and dimensionality of the code, respectively.
@@
The encoder comprises parallel amplitude and phase sub-encoders that share an identical network architecture, as shown in Fig. 1. In the amplitude/phase sub-encoder, the input amplitude/phase spectrum is initially processed through an input 1D convolutional layer (channel size $=C$), a layer normalization [34] and then undergoes deep processing through a modified ConvNeXt v2 network. The output of the modified ConvNeXt v2 network is further processed by a layer normalization and a feed-forward layer with $C$ nodes. The 1D downsampling convolutional layer (channel size $=C/2$ and stride $=D$) serves as the ultimate component in the amplitude/phase sub-encoder, further downsampling the output features of the feed-forward layer by a factor of $D$ and reducing its dimensionality by half to generate amplitude/phase continuous latent code $\bm{C}_{A}$/$\bm{C}_{P}\in\mathbb{R}^{F_{c}\times(C/2)}$. Therefore, we have $F_{c}=F/D$.
@@
The modified ConvNeXt v2 network, which is constructed by cascading 8 identical modified ConvNeXt v2 blocks, serves as the backbone of both encoder and decoder. The modified ConvNeXt v2 block has been adapted from the ConvNeXt v2 block originally designed for image processing [35]. The primary modification involves replacing 2D convolutions with 1D ones, thereby tailoring the block to better suit the processing of audio signals. As shown in Fig. 2, in each modified ConvNeXt v2 block, the input sequentially passes through a 1D depth-wise convolutional layer (channel size $=C$), a layer normalization, a feed-forward layer with $C_{H}$ nodes that projects features into a higher dimensionality (i.e., $C_{H}>C$), a Gaussian error linear unit (GELU) activation [36], a global response normalization (GRN) layer [35], an another feed-forward layer with $C$ nodes that projects features back to the original dimensionality, and finally superimposes with the input (i.e., residual connections) to obtain the output.
@@
To aggregate both the amplitude and phase information, we concatenate the amplitude code and phase code along the dimension axis to obtain a fused latent code $\left[\bm{C}_{A},\bm{C}_{P}\right]\in\mathbb{R}^{F_{c}\times C}$. Then, a dimensionality-reduction 1D convolutional layer (channel size $=N_{c}$) is used to significantly reduce the dimension of this fused code, resulting in a low-dimensional fused continuous latent code $\bm{C}\in\mathbb{R}^{F_{c}\times N_{c}}$ which combines both the amplitude and phase information. The reason for reducing the dimensionality of the continuous latent code $\bm{C}$ is to concurrently decrease the dimensionality of codebooks during subsequent quantization process, facilitating the storage and transmission of codebooks. The sampling rate of $\bm{C}$ is $f_{s}/w_{s}/D$, which is one-$D$ of the sampling rate of the amplitude and phase spectra.
@@ III-A2 Quantizer
III-A3 Decoder
As illustrated in Fig. 1, the decoder decodes the log amplitude spectrum $\hat{\bm{A}}\in\mathbb{R}^{F\times N}$ and phase spectrum $\hat{\bm{P}}\in\mathbb{R}^{F\times N}$ in parallel from the input quantized latent code $\hat{\bm{C}}\in\mathbb{R}^{F_{c}\times N_{c}}$, and finally reconstructs the decoded waveform $\hat{\bm{x}}\in\mathbb{R}^{T}$ through ISTFT. The structure of the decoder is roughly symmetrical to that of the encoder. The parallel amplitude and phase sub-decoders are the primary components of the decoder. The quantized latent code $\hat{\bm{C}}$ is first dimensionally restored through a 1D dimensionality-augmentation convolutional layer (channel size $=C/2$) to be used as input for both the amplitude and phase sub-decoders. In the amplitude sub-decoder, the input is first upsampled $D$ times through a 1D deconvolutional layer (channel size $=C$ and stride $=D$), a layer normalization and then processed by a modified ConvNeXt v2 network along with a layer normalization and a feed-forward layer with $C$ nodes. Finally, a 1D output convolutional layer (channel size $=N$) is adopted to predict the decoded log amplitude spectrum $\hat{\bm{A}}$. The sole distinction between the phase sub-decoder and the amplitude sub-decoder lies in the utilization of a phase parallel estimation architecture proposed in our previous publication [37] at the output end of the phase sub-decoder. The parallel estimation architecture ensures the direct prediction of wrapped phase spectra, consisting of two identical parallel 1D convolutional layers (channel size $=N$) and a phase calculation formula $\bm{\Phi}$. Assume the outputs of the two parallel layers are $\hat{\bm{R}}\in\mathbb{R}^{F\times N}$ and $\hat{\bm{I}}\in\mathbb{R}^{F\times N}$, respectively, then the phase spectrum is calculated by $\hat{\bm{P}}=\bm{\Phi}(\hat{\bm{R}},\hat{\bm{I}})$. Function $\bm{\Phi}$ is calculated element-wise. For $\forall R\in\mathbb{R}$ and $I\in\mathbb{R}$, we have

### v2

@@ Abstract
This paper introduces a novel neural audio codec targeting high waveform sampling rates and low bitrates named APCodec, which seamlessly integrates the strengths of parametric codecs and waveform codecs. The APCodec revolutionizes the process of audio encoding and decoding by concurrently handling the amplitude and phase spectra as audio parametric characteristics like parametric codecs. It is composed of an encoder and a decoder with the modified ConvNeXt v2 network as the backbone, connected by a quantizer based on the residual vector quantization (RVQ) mechanism. The encoder compresses the audio amplitude and phase spectra in parallel, amalgamating them into a continuous latent code at a reduced temporal resolution. This code is subsequently quantized by the quantizer. Ultimately, the decoder reconstructs the audio amplitude and phase spectra in parallel, and the decoded waveform is obtained by inverse short-time Fourier transform. To ensure the fidelity of decoded audio like waveform codecs, spectral-level loss, quantization loss, and generative adversarial network (GAN) based loss are collectively employed for training the APCodec. To support low-latency streamable inference, we employ feed-forward layers and causal deconvolutional layers in APCodec, incorporating a knowledge distillation training strategy to enhance the quality of decoded audio. Experimental results confirm that our proposed APCodec can encode 48 kHz audio at bitrate of just 6 kbps, with no significant degradation in the quality of the decoded audio. At the same bitrate, our proposed APCodec also demonstrates superior decoded audio quality and faster generation speed compared to well-known codecs, such as Encodec, AudioDec and DAC.
@@ I Introduction
The waveform codecs aim at encoding the input audio waveform directly and reproducing a faithful reconstruction of the input audio waveform, such as pulse code modulation (PCM) [20]. Although waveform codecs can decode high-quality audio, they also require a higher bitrate, which increases storage and transmission costs. In recent times, end-to-end neural waveform codecs with raw waveform I/O have surfaced, offering a partial equilibrium between decoded audio quality and bitrate [21, 22, 23, 24, 25, 26]. For example, SoundStream [25] and Encodec [26] employed the residual vector quantization (RVQ) mechanism [27] to reduce the bitrate, while utilizing the losses from the HiFi-GAN vocoder [28] to ensure the fidelity of the decoded audio. Recently, researchers have made improvements addressing the issues present in current end-to-end audio codecs, primarily focusing on quantization strategies. On the one hand, in applications such as audio communication, audio codecs have incorporated variations of RVQ to decrease bitrates and improve communication efficiency [29, 30, 31]. For example, HiFi-Codec [29] has introduced group RVQ (GRVQ) to reduce information redundancy in RVQ. This allows for high-quality audio coding with less codebooks, resulting in a reduced bitrate. On the other hand, in downstream tasks such as TTS, efforts have been undertaken to introduce or disentangle semantic information during the quantization stage, aiming to tailor the approach to the specific tasks [7, 8, 9]. Moreover, there have been endeavors to improve the model structure [30] or incorporate additional signal processing techniques (e.g., bandwidth extension) in codecs [32]. Although these codecs have indeed enhanced the decoded audio quality and decreased bitrates, they still required more than a hundred times the downsampling and upsampling operations due to the direct waveform encoding and decoding, leading to high model complexity. Besides, direct waveform encoding and decoding could also potentially result in low generation efficiency. Some recent works have also overlooked considerations for low latency, making it challenging to achieve streamable inference [29, 30].
@@
Beyond the aforementioned challenges, encompassing bitrate, generation speed, and latency in existing audio codecs, there is presently scant research devoted to audio codecs tailored for higher waveform sampling rates (e.g., 48 kHz). Currently, neural audio codecs (e.g., SoundStream [25] and HiFi-Codec [29]) are mostly designed for processing audio at sampling rates of 16 kHz or 24 kHz. This limitation hinders the utilization of audio codecs for compressing high-sampling-rate audio data and poses challenges for downstream tasks like TTS, which aim to meet the demand for higher quality speech generation. The aforementioned MDCT-based parametric codecs [18, 19], while targeted at 48 kHz audio, demands an excessively high bitrate. Although AudioDec [33] can achieve 48 kHz audio codec at bitrate of 12.8 kbps, it still requires the integration of a neural vocoder and the adoption of a multi-stage training strategy, as reported in [33]. Descript audio codec (DAC) [34] can achieve 44.1 kHz audio coding at a bitrate of only 8 kbps, thanks to improved RVQ for improving codebook utilization and improved losses to enhance the decoded audio quality. However, the DAC’s bitrate remains relatively high and lacks consideration for low latency.
@@
To address the aforementioned challenges, this paper proposes a novel neural audio codec named APCodec. It endeavors to provide high-quality decoded audio, while maintaining a low bitrate, fast generation speed, and low latency, specifically tailored for 48 kHz audio. Like parametric codecs, the proposed APCodec regards amplitude and phase spectra as audio parametric characteristics during the encoding and decoding processes, rather than directly processing the raw waveform. A notable advantage of this approach lies in its simplicity, as it only requires uncomplicated downsampling to obtain latent codes at an appropriately low sampling rate, thereby effectively reducing the bitrate. RVQ [27] is also utilized for code quantization to further reduce the bitrate. With the objective of achieving faithful waveform reconstruction akin to waveform codecs, a comprehensive combination of spectral-level loss, quantization loss and generative adversarial network (GAN) based loss are employed to train the APCodec. To attain streamable inference, a low-latency implementation is achieved by integrating feed-forward layers and causal deconvolutional layers, complemented by the application of a knowledge distillation training strategy. The resulting fixed latency is only 6.67 ms for the 48 kHz audio codec. Experimental results have confirmed that the proposed APCodec can achieve high-quality 48 kHz audio coding at a bitrate of only 6 kbps with only 8$\times$ downsampling/upsampling. At the same bitrate, our proposed APCodec significantly outperforms several well-known neural codecs which support high-sampling-rate audio coding, e.g., Encodec [26], AudioDec [33] and DAC [34] in terms of decoded audio quality. The APCodec also demonstrates the fastest generation speed, attaining an impressive 89$\times$ real-time performance on GPU and 5.8$\times$ real-time performance on CPU. This remarkable acceleration is attributed to its comprehensive all-frame-level processing.
@@
This paper is organized as follows: In Section II, we provide details on our proposed APCodec. In Section III, we present our experimental results. Finally, we give conclusions in Section IV.
Figure 1: Details of the model structure of the proposed APCodec. Here, Conv1D, DeConv1D, Concat, $\Phi$, STFT and ISTFT represent the 1D convolutional layer, 1D deconvolutional layer, concatenation, phase calculation formula, short-time Fourier transform and inverse short-time Fourier transform, respectively. For waveforms, the content after @ represents the sampling rate, while for spectra and codes, the content after @ represents the frame rate (taking a sampling rate of 48 kHz and a bitrate of 6 kbps as an example).
II Proposed Methods
Unlike some well-known neural waveform codecs, e.g., SoundStream [25], Encodec [26], HiFi-Codec [29], AudioDec [33] and DAC [34], our proposed APCodec encodes and quantizes amplitude and phase spectra extracted from the audio waveform through short-time Fourier transform (STFT). Finally, it decodes the quantized codes into amplitude and phase spectra and restores the audio waveform through inverse STFT (ISTFT). Subsequently, we will present a detailed overview of the model structure and training criteria of the proposed APCodec. Additionally, we will discuss the low-latency implementation for APCodec.
II-A Model Structure
@@ II-A1 Encoder
As illustrated in Fig. 1, the encoder takes the log amplitude spectrum $\bm{A}\in\mathbb{R}^{F\times N}$ and phase spectrum $\bm{P}\in\mathbb{R}^{F\times N}$ extracted from the audio waveform $\bm{x}\in\mathbb{R}^{T}$ using STFT as inputs and encodes them into a continuous latent code $\bm{C}\in\mathbb{R}^{F_{c}\times N_{c}}$ that contains fused amplitude and phase information in parallel. Here, $T$ represents the number of time-domain waveform samples, and $F$ and $N$ respectively represent the number of spectral frames and frequency bins. Assuming the sampling rate of $\bm{x}$ is $f_{s}$ and the frame shift of the STFT is $w_{s}$, the resulting frame rate for the extracted amplitude and phase spectra is $f_{s}/w_{s}$, and it holds that $T=F\cdot w_{s}$. $F_{c}$ and $N_{c}$ denote the number of frames and dimensionality of the code, respectively.
@@
The encoder comprises parallel amplitude and phase sub-encoders that share an identical network architecture, as shown in Fig. 1. In the amplitude/phase sub-encoder, the input log amplitude/phase spectrum is initially processed through an input 1D convolutional layer (channel size $=K$), a layer normalization [35] and then undergoes deep processing through a modified ConvNeXt v2 network. The output of the modified ConvNeXt v2 network is further processed by a layer normalization and a feed-forward layer with $K$ nodes. The 1D downsampling convolutional layer (channel size $=K/2$ and stride $=D$) serves as the ultimate component in the amplitude/phase sub-encoder, further downsampling the output features of the feed-forward layer by a factor of $D$ and reducing its dimensionality by half to generate amplitude/phase continuous latent code $\bm{C}_{A}$/$\bm{C}_{P}\in\mathbb{R}^{F_{c}\times(K/2)}$. Therefore, we have $F_{c}=F/D$.
@@
The modified ConvNeXt v2 network, which is constructed by cascading 8 identical modified ConvNeXt v2 blocks, serves as the backbone of both encoder and decoder. The modified ConvNeXt v2 block has been adapted from the ConvNeXt v2 block originally designed for image processing [36]. The primary modification involves replacing 2D convolutions with 1D ones, thereby tailoring the block to better suit the processing of audio signals. As shown in Fig. 2, in each modified ConvNeXt v2 block, the input sequentially passes through a 1D depth-wise convolutional layer (channel size $=K$), a layer normalization, a feed-forward layer with $K_{H}$ nodes that projects features into a higher dimensionality (i.e., $K_{H}>K$), a Gaussian error linear unit (GELU) activation [37], a global response normalization (GRN) layer [36], an another feed-forward layer with $K$ nodes that projects features back to the original dimensionality, and finally superimposes with the input (i.e., residual connections) to obtain the output.
@@
To aggregate both the amplitude and phase information, we concatenate the amplitude code and phase code along the dimension axis to obtain a fused latent code $\left[\bm{C}_{A},\bm{C}_{P}\right]\in\mathbb{R}^{F_{c}\times K}$. Then, a dimensionality-reduction 1D convolutional layer (channel size $=N_{c}$) is used to significantly reduce the dimension of this fused code, resulting in a low-dimensional fused continuous latent code $\bm{C}\in\mathbb{R}^{F_{c}\times N_{c}}$ which combines both the amplitude and phase information. The reason for reducing the dimensionality of the continuous latent code $\bm{C}$ is to concurrently decrease the dimensionality of codebooks during subsequent quantization process, facilitating the storage and transmission of codebooks. The frame rate of $\bm{C}$ is $f_{s}/w_{s}/D$, which is one-$D$ of the frame rate of the amplitude and phase spectra.
@@ II-A2 Quantizer
II-A3 Decoder
As illustrated in Fig. 1, the decoder decodes the log amplitude spectrum $\hat{\bm{A}}\in\mathbb{R}^{F\times N}$ and phase spectrum $\hat{\bm{P}}\in\mathbb{R}^{F\times N}$ in parallel from the input quantized latent code $\hat{\bm{C}}\in\mathbb{R}^{F_{c}\times N_{c}}$, and finally reconstructs the decoded waveform $\hat{\bm{x}}\in\mathbb{R}^{T}$ through ISTFT. The structure of the decoder is roughly symmetrical to that of the encoder. The parallel amplitude and phase sub-decoders are the primary components of the decoder. The quantized latent code $\hat{\bm{C}}$ is first dimensionally restored through a 1D dimensionality-augmentation convolutional layer (channel size $=K/2$) to be used as input for both the amplitude and phase sub-decoders. In the amplitude sub-decoder, the input is first upsampled $D$ times through a 1D deconvolutional layer (channel size $=K$ and stride $=D$), a layer normalization and then processed by a modified ConvNeXt v2 network along with a layer normalization and a feed-forward layer with $S$ nodes. Finally, a 1D output convolutional layer (channel size $=N$) is adopted to predict the decoded log amplitude spectrum $\hat{\bm{A}}$. The sole distinction between the phase sub-decoder and the amplitude sub-decoder lies in the utilization of a phase parallel estimation architecture proposed in our previous publication [38] at the output end of the phase sub-decoder. The parallel estimation architecture ensures the direct prediction of wrapped phase spectra, consisting of two identical parallel 1D convolutional layers (channel size $=N$) and a phase calculation formula $\bm{\Phi}$. Assume the outputs of the two parallel layers are $\hat{\bm{R}}\in\mathbb{R}^{F\times N}$ and $\hat{\bm{I}}\in\mathbb{R}^{F\times N}$, respectively, then the phase spectrum is calculated by $\hat{\bm{P}}=\bm{\Phi}(\hat{\bm{R}},\hat{\bm{I}})$. Function $\bm{\Phi}$ is calculated element-wise. For $\forall R\in\mathbb{R}$ and $I\in\mathbb{R}$, we have